# SEA: Self-Evolution Agent with Step-wise Reward for Computer Use

### v1

@@ Abstract
Computer use agent is an emerging area in artificial intelligence that aims to operate the computers to achieve the user’s tasks, which attracts a lot of attention from both industry and academia. However, the present agents’ performance is far from being used. In this paper, we propose the Self-Evolution Agent (SEA) for computer use, and to develop this agent, we propose creative methods in data generation, reinforcement learning, and model enhancement. Specifically, we first propose an automatic pipeline to generate the verifiable trajectory for training. And then, we propose efficient step-wise reinforcement learning to alleviate the significant computational requirements for long-horizon training. In the end, we propose the enhancement method to merge the grounding and planning ability into one model without any extra training. Accordingly, based on our proposed innovation of data generation, training strategy, and enhancement, we get the Self-evolution Agent (SEA) for computer use with only 7B parameters, which outperforms models with the same number of parameters and has comparable performance to larger ones. We will make the models’ weight and related codes open-source in the future.
footnotetext: Corresponding author: Yukang Huo. Email: yukanghuo.ai@gmail.com
I Introduction
@@
Computer use agent has been an attractive technology that follows the users’ instructions to manipulate computers and match the users’ expectations, and it has been a milestone towards general artificial intelligence. In order to achieve the goal of computer use, the agent should understand the computer environment, provide a plan based on the instructions, and finally manipulate the computer. For the computer environment, researchers will take a screenshot as a representation of that. With the development of the MLLMs [1, 2], they have the remarkable ability of perception and reasoning, which can meet the requirements of a computer-use agent.
Figure 1: The comparison with the SOTA methods
@@
Despite significant progress in the field of intelligent agents, building robust agents for real-world deployment remains highly challenging, particularly in scenarios that involve complex multimodal reasoning and dynamic, uncertain environments. One of the most critical challenges in training intelligent agents is acquiring high-quality data. Current mainstream data construction approaches predominantly rely on manually designed task templates and annotations. This process is not only costly and difficult to scale but also often fails to meet the semantic coverage and diversity required by complex tasks. Moreover, real-world tasks typically involve multi-step, long-horizon interactions, where informative reward signals only emerge after completing the entire task or passing through several key steps. Such sparse reward settings make it difficult for the agent to establish a clear correlation between individual actions and final outcomes, leading to suboptimal task performance. However, when handling complex GUI tasks, encoding image data often requires processing a large sequence of tokens, many of which contribute minimally to task execution and reasoning but incur significant computational costs. In addition, existing reasoning mechanisms frequently depend on external models to assist in evaluating and selecting trajectories. This reliance introduces significant computational overhead and increases system complexity, further limiting the efficiency and scalability of the overall agent architecture.
@@
To address these challenges, we propose a set of novel agent training methods, designed to enable efficient and scalable task learning and execution. In the data acquisition phase, we propose a closed-loop pipeline to generate the verifiable tasks, which leverages a task-generation agent and a code-generation agent to automatically synthesize tasks along with their corresponding execution and verification Python programs. This dual-code generation strategy ensures that each task is accompanied by a well-defined success criterion and a programmatic validation method, effectively forming a closed-loop supervision structure. By tightly coupling the task definition with its success condition, the proposed pipeline significantly enhances both the effectiveness and usability of the generated data. Furthermore, we use virtual environment to run each task using its corresponding execution program and applying the verification program to retain only those tasks that are successfully completed. And then we propose the Generation and Assessment for Trajectory Extraction (GATE) method, designed to construct high-quality and representative training trajectories that improve agent learning efficiency. Specifically, for each validated task, we perform multiple rounds of inference to generate diverse candidate trajectories. As training progresses, we gradually replace the inference model with our agent during the middle and later stages of training. Each generated trajectory is then filtered again using the verification code to ensure successful task completion. Among the successful trajectories for the same task, we prioritize those with the fewest required steps, selecting more efficient execution paths. Additionally, we train a step filtering model to eliminate redundant or incorrect steps within each trajectory, resulting in high-quality, concise training data.
@@
During the training phase, we introduce the Trajectory Reasoning by Step-wise Reinforcement Learning(TR-SRL). This component evaluates each step in a task trajectory by comparing the agent’s execution results against the corresponding ground-truth outcomes. In addition, TR-SRL assesses the consistency between the agent’s generated thoughts and executed actions at each step, ensuring alignment between internal reasoning and external behavior.
@@
At last, we propose the method to further enhance the generalization ability and effect of the model. Specifically, we first train a grounding model and integrate its weights, which effectively improves the performance of both grounding and planning. In addition, we introduce a Temporal Compressed Sensing Mechanism to optimize perception efficiency by modeling the importance of input image tokens over time. By retaining only the most important tokens, the model achieves a significant reduction in training cost while maintaining critical semantic content and improving inference efficiency.
@@
In summary, based on these proposed methods including verifiable data generation, efficient reinforcement learning, and model enhancement, we create the SEA, which is the Self-Evolution Agent for computer use. As shown in the 1, our agent outperforms the 7B models and has a comparable performance with a larger number of parameters.
II Related Works
II-A High-quality Training Data
@@
High-quality data is essential for bridging the gap between the symbolic world [3] and the digital world [4], thereby advancing the development of intelligent agents. Early works such as Rico [5] and Mini-Wob [6] provided serialized GUI data and low-level keyboard/mouse interactions for mobile applications and web-based tasks, respectively. Subsequent studies have further expanded data resources for mobile [7, 8, 9, 10], web [11, 12, 13], and desktop applications [14]. To train agents more efficiently, researchers typically use trajectory data, which consists of sequences containing GUI information, including both low-level and high-level instructions and corresponding operations [15, 16, 17]. In addition, some studies utilize demonstration data, filtered imitation learning, and self-collected data [18, 19] to fine-tune vision-language models (VLMs), such as AutoUI and CogAgent [20, 21]. PAE [22] generates tasks automatically to provide diverse training data for agents.
@@
However, these methods face several challenges in practical applications: the reliance on costly manually annotated task templates, the inability of static data to adapt to real-world dynamics, and the inconsistent quality of task trajectories—all of which severely limit the effectiveness and adaptability of agents in diverse environments. So in our paper, we have proposed a pipeline to generate high-quality verifiable tasks with correct trajectory.
II-B Compute Use Agent
Recent advancements in multi-modal large language models (MLLMs) have enabled agents to perform increasingly complex tasks within graphical user interfaces (GUIs). Models such as SeeClick [23] and ScreenAgent [24] leverage vision-language models (VLMs) to interpret UI screens and execute interactive tasks based on multimodal prompts. GUI-R1 [25] enhances UI grounding performance by introducing reinforcement learning mechanisms. Song [26] proposes a framework for API-driven web agents that utilizes task-specific background knowledge to execute complex web operations. The AgentWorkflow Memory (AWM) module [27] further optimizes memory management by selectively providing relevant workflows to guide the agent’s subsequent actions. Meanwhile, there are some specialized computer-use agents, such as PresentAgent [28], that aim to help users generate the presentation. Despite the substantial progress made in perception, understanding, and execution, existing approaches still face key limitations: sparse rewards in long-horizon tasks, high computational costs during training, and a lack of modeling for the consistency between thought and action within action sequences. So in this paper, we propose to use the step-wise RL training to replace the long-horizon training, which is achieved by the dedicated designing of rewards and the advanced RL training strategy.
II-C Reinforcement Learning
@@
As the fascinating performance of the DeepSeek-R1 [29], researchers have noticed that the great of the Post-training in improving the performance of the LLM, especially comparing with the minor imporvence of the pre-trainig phase. OpenAI’s o1 [30] and DeepSeek-R1 [29] demonstrate strong capabilities in tasks such as mathematical reasoning [31], code generation [32], and multimodal reasoning [33, 34] through structured reward signals. Sweet-RL [35] introduces a multi-round DPO framework to improve long-horizon behavior in language agents, while ARPO [36] increases success rates on complex multi-turn GUI tasks by incorporating experience replay and task selection strategies. DISTRL [37] boosts training efficiency and success rates for GUI agents in mobile device control tasks through its novel asynchronous distributed reinforcement learning framework and the custom A-RIDE algorithm, while also enhancing generalization across diverse tasks. UI-TARS [38] significantly improves performance in complex GUI tasks through innovations in enhanced perception, unified action modeling, and dual-stage reasoning. Moreover, some methods propose interactive refinement of LLMs/VLMs—particularly Web/GUI agents—using self-evaluated feedback [18, 39, 40, 37].
III Method
In this section, we provide a detailed description of the proposed method. Firstly, we will provide a simple review of reinforcement learning. We will then dive into our proposed method. As we have mentioned, the proposed method consists of three parts: automatic data collection, RL with multiple rewards, and model enhancement with a grounding model.
III-A Data Engine for Self-evolution
One of the key points to make a model evolve is that we could have the ability to scale up the training data, which will provide the model more probability to explore. Nevertheless, the collection of this kind of data, operating the computer in a human-style, is so complex that the amount of this data is insufficient for scaling up. To solve this challenge, we propose a novel method to generate a trajectory dataset of manipulating different computer software. Specifically, the proposed method can be divided into three parts: (1)Generation of closed-loop verifiable tasks; (2) Generation and Assessment for Trajectory Extract; (3) Filtering of data with a step model.
III-A1 Generation of closed-loop verifiable tasks
@@
To enable the agent to better handle complex multimodal tasks and dynamic environment interactions, we design a closed-loop task data generation pipeline, as illustrated in Figure 2. Firstly, the proposed pipeline will use a Task Agent to produce many task instructions. The task agent will use some actual tasks of a specific software as the few-shot examples to evoke the agent’s in-context learning of the target sofware. During the generation process, the task agent will periodically check whether the generated tasks can be executed and verified and whether there are some similar tasks have been generated. Meanwhile, the agent will produce some guidelines of how to complete the task. Secondly, the code generation agent will take the generated task instruction and the guideline as inputs to generate a batch of execution and verification program. For execution program, it’s target is to achieve tasks’ goals described in instructions by using the Python. And for the verification program, it judges whether the execution programs’ output matches the tasks’ requirement. In the end, we will automatically run pairs of execution program and verification to select the correct verification program. So in this way, we can get the verifiable tasks.
Figure 2: Illustration of Generating Closed-loop Task Data.
Figure 3: Illustration of the data generation and multi-stage trajectory filtering strategy.
III-A2 Generation and Assessment for Trajectory Extraction
@@
To enable the agent to learn the most relevant and effective action sequences, we propose a novel multi-stage trajectory filtering strategy, as illustrated in Figure 3. After getting the task instruction and verification program, we will firstly use our cold-start model, UI-TARS-1.5-7B [38], to perform multiple rounds of inference and obtain diverse task trajectories. And the verification program is then used to select only those trajectories that lead to successful task completion. For each task, we prioritize trajectories with the minimal number of steps based on step-length information. Additionally, to further ensure the quality of each step within the trajectory, we train a step filtering model to filter out redundant or erroneous steps, thereby yielding optimal training trajectories. And the step model will be introduced in the next section. It is notable that during the training process, we will gradually replace the cold-start model with our trained model to generate trajectories for later training stages. In this way, we can make our model evolve through the whole training phase.
III-B Filtering data with step model

### v2

# Evolving in Tasks: Empowering the Multi-modality Large Language Model as the Computer Use Agent

@@ Abstract.
Computer use agents represent an emerging area in artificial intelligence, aiming to operate computers autonomously to fulfill user tasks, attracting significant attention from both industry and academia. However, the performance of existing agents remains insufficient for practical deployment. In this paper, we propose the Self-Evolution Agent (SEA) for computer operation, alongside three core innovations in data generation, reinforcement learning, and model enhancement to develop this agent. Specifically, we first design an automatic pipeline to generate verifiable task trajectories for training. Second, we propose Efficient Step-wise Reinforcement Learning to reduce the substantial computational overhead of long-horizon training. Finally, we introduce a model enhancement method that integrates grounding and planning capabilities into a single model without additional training. Leveraging these innovations, our SEA (with only 7B parameters) outperforms existing models of the same parameter scale and achieves performance comparable to larger models (e.g., 32B/72B parameters) on computer use tasks. We plan to release the model weights and related code as open-source resources in the future.
Browser and Computer Use Agent; Reinforcement Learning; Multi-modality Large Language Model
1. Introduction
@@
Computer-use agents have been an attractive technology that follows users’ instructions to manipulate computers and match their expectations, marking a milestone towards general artificial intelligence. As the browser is the most important software on a computer, the agent’s functions are divided into two categories. The first one involves manipulating the web browser, and the second consists of using other software, such as office applications, media players, IDEs, and so on. To achieve the goal of computer use, the agent must understand the computer environment, provide a plan based on the instructions, and ultimately manipulate the computer. For the computer environment, researchers will take a screenshot as a representation of that. With the development of Multi-modal Large Language Models (MLLMs), such as (Shuai et al., 2025; Guo et al., 2025), they possess a remarkable ability for perception and reasoning, which can meet the requirements of a computer-use agent. Meanwhile, these models satisfy the requirement of making execution plans and operating the computer after some post-training.
Figure 1. The comparison with the SOTA methods. This scatter plot depicts the OSWorld success rates of different open-source and closed-source models. Our proposed model achieves the highest success rate among open-source 7B models and substantially outperforms some other open-source alternatives with a larger number of parameters, even coming close to the performance of certain closed-source models.
@@
Despite significant progress in the field of intelligent agents, building robust agents for real-world deployment remains highly challenging, particularly in scenarios that involve complex multimodal reasoning and dynamic, uncertain environments. For example, when operating websites, there are always some pop-up advertisements, and their contents vary from time to time. One of the most critical challenges in training intelligent agents is acquiring high-quality data. Current mainstream data construction approaches predominantly rely on manually designed task templates and annotations. This process is not only costly and difficult to scale but also often fails to meet the semantic coverage and diversity required by complex tasks. Moreover, real-world tasks typically involve multi-step, long-horizon interactions, where informative reward signals only emerge after completing the entire task or passing through several key steps. Such sparse reward settings make it difficult for the agent to establish a clear correlation between individual actions and final outcomes, leading to suboptimal task performance. However, when handling complex GUI tasks, encoding image data often requires processing a large sequence of tokens, many of which contribute minimally to task execution and reasoning but incur significant computational costs. In addition, existing reasoning mechanisms frequently depend on external models to assist in evaluating and selecting trajectories. This reliance introduces significant computational overhead and increases system complexity, further limiting the efficiency and scalability of the overall agent architecture.
@@
To address these challenges, we propose a set of novel agent training methods, designed to enable efficient and scalable task learning and execution. In the data acquisition phase, we propose a closed-loop pipeline to generate the verifiable tasks, which leverages a task-generation agent and a code-generation agent to automatically synthesize tasks along with their corresponding execution and verification Python programs. This dual-code generation strategy ensures that each task is accompanied by a well-defined success criterion and a programmatic validation method, effectively forming a closed-loop supervision structure. By tightly coupling the task definition with its success condition, the proposed pipeline significantly enhances both the effectiveness and usability of the generated data. Furthermore, we use a virtual environment to run each task using its corresponding execution program and apply the verification program to retain only those tasks that are successfully completed.
And then we propose the Generation and Assessment for Trajectory Extraction (GATE) method, designed to construct high-quality and representative training trajectories that improve agent learning efficiency. Specifically, for each validated task, we perform multiple rounds of inference to generate diverse candidate trajectories. As training progresses, we gradually replace the inference model with our agent during the middle and later stages of training. Each generated trajectory is then filtered again using the verification code to ensure successful task completion. Among the successful trajectories for the same task, we prioritize those with the fewest required steps, selecting more efficient execution paths. Additionally, we train a step filtering model to eliminate redundant or incorrect steps within each trajectory, resulting in high-quality, concise training data.
@@
Another challenge is how to efficiently use the sparse rewards, so during the training phase, we introduce the Trajectory Reasoning by Step-wise Reinforcement Learning(TR-SRL). This component evaluates each step in a task trajectory by comparing the agent’s execution results against the corresponding ground-truth outcomes. In addition, TR-SRL assesses the consistency between the agent’s generated thoughts and executed actions at each step, ensuring alignment between internal reasoning and external behavior.
@@
At last, we propose a grounding-based enhancement method to further improve the generalization ability and effect of the model. Specifically, we first train a grounding model and integrate its weights, which effectively improves the performance of both grounding and planning. In addition, we introduce a Temporal Compressed Sensing Mechanism to optimize perception efficiency by modeling the importance of input image tokens over time. By retaining only the most essential tokens, the model achieves a significant reduction in training cost while maintaining critical semantic content and improving inference efficiency.
@@
In summary, based on these proposed methods, including verifiable data generation, efficient reinforcement learning, and model enhancement, we create the Self-Evolution Agent (SEA) for computer use. This agent utilizes the synthesis of verifiable tasks to evolve through step-wise reinforcement learning. As shown in  Figure 1, our agent outperforms the 7B models and achieves comparable performance despite having a larger number of parameters.
2. Related Works
2.1. High-quality Training Data
@@
High-quality data is essential for bridging the gap between the symbolic world (Fangzhi et al., 2023) and the digital world (Zhiyong et al., 2024a), thereby advancing the development of intelligent agents. Early works such as Rico (Biplab et al., 2017) and Mini-Wob (Tianlin et al., 2017) provided serialized GUI data and low-level keyboard/mouse interactions for mobile applications and web-based tasks, respectively. Subsequent studies have further expanded data resources for mobile (Christopher et al., 2023; Jiwen et al., 2024; Quanfeng et al., 2024; Yuxiang et al., 2024), web (Zheran et al., 2018; Lù et al., 2024; Murty et al., 2024), and desktop applications (Chen et al., 2024). To train agents more efficiently, researchers typically use trajectory data, which consists of sequences containing GUI information, including both low-level and high-level instructions and corresponding operations (Wei et al., 2024; Jianguo et al., 2024; Longtao et al., 2024). In addition, some studies utilize demonstration data, filtered imitation learning, and self-collected data (Jiayi et al., 2024; Lai et al., 2024) to fine-tune vision-language models (VLMs), such as AutoUI and CogAgent (Raghav et al., 2024; Zhang and Zhang, 2023). PAE (Yifei et al., 2025a) generates tasks automatically to provide diverse training data for agents.
@@
However, these methods face several challenges in practical applications: reliance on costly manually annotated task templates, inability of static data to adapt to real-world dynamics, and inconsistent quality of task trajectories—all of which severely limit the effectiveness and adaptability of agents in diverse environments. To address these limitations, in this paper, we propose a closed-loop pipeline to generate high-quality, verifiable tasks with correct execution trajectories, ensuring both data scalability and reliability.
2.2. Browser and Computer Use Agent
Recent advancements in multi-modal large language models (MLLMs) have empowered agents to tackle increasingly complex tasks within graphical user interfaces (GUIs). Models such as SeeClick (Kanzhi et al., 2024) and ScreenAgent (Runliang et al., 2024) utilize vision-language models (VLMs) for interpreting UI screens and executing interactive tasks based on multimodal prompts. GUI-R1 (Run et al., 2025) further enhances UI grounding performance by integrating reinforcement learning (RL) mechanisms into the agent architecture.
For web-centric scenarios, Song (Song et al., 2024) proposes an API-driven web agent framework that incorporates task-specific background knowledge to execute complex network operations. Distinct from API-reliant counterparts, modern browser agents emphasize human-like interactive capabilities: WebSailor (Li et al., 2025b), an open-source browser agent, combines multi-modal web perception (HTML structure + visual elements) with the DUPO RL algorithm to handle cross-page tasks and dynamic content loading, achieving state-of-the-art performance on benchmarks like BrowseComp. Complementing these, the AgentWorkflow Memory (AWM) module (Wang et al., 2024b) optimizes memory management by selectively retrieving task-relevant workflows to guide sequential actions. Beyond general-purpose agents, specialized tools like PresentAgent (Shi et al., 2025) have emerged, which specializes in automated presentation generation.
Despite substantial progress in perception, reasoning, and execution, existing approaches—particularly browser agents—face pronounced limitations: (1) Sparse rewards are exacerbated in long-horizon browser tasks (e.g., multi-step e-commerce operations) due to intermittent feedback from dynamic web content; (2) Training costs remain prohibitive for agents handling diverse browser environments (e.g., varying UI layouts and JavaScript rendering); (3) Few methods model the thought-action consistency in action sequences, leading to misalignment between planned operations (e.g., ”click checkout”) and actual GUI interactions. To address these issues, this paper proposes a step-wise RL training paradigm, replacing monolithic long-horizon training through tailored reward design and advanced RL strategy optimization.
2.3. Reinforcement Learning
@@
As the fascinating performance of the DeepSeek-R1 (Daya et al., 2025), researchers have noticed that the great improvement of the Post-training phase in improving the performance of the LLM, especially compared to the minor improvement of the pre-training phase. OpenAI’s o1 (Aaron et al., 2024) and DeepSeek-R1 (Daya et al., 2025) demonstrate strong capabilities in tasks such as mathematical reasoning (Zhihong et al., ), code generation (Liu and Zhang, 2025), and multimodal reasoning (Wenxuan et al., 2025; Yuqi et al., 2025) through structured reward signals. Sweet-RL (Yifei et al., 2025b) introduces a multi-round DPO framework to improve long-horizon behavior in language agents, while ARPO (Lu et al., 2025) increases success rates on complex multi-turn GUI tasks by incorporating experience replay and task selection strategies. DISTRL (Taiyi et al., 2024) boosts training efficiency and success rates for GUI agents in mobile device control tasks through its novel asynchronous distributed reinforcement learning framework and the custom A-RIDE algorithm, while also enhancing generalization across diverse tasks. UI-TARS (Yujia et al., 2025) significantly improves performance in complex GUI tasks through innovations in enhanced perception, unified action modeling, and dual-stage reasoning. Moreover, some methods propose interactive refinement of LLMs/VLMs—particularly Web/GUI agents—using self-evaluated feedback (Jiayi et al., 2024; Hao et al., 2024; Pranav et al., 2024; Taiyi et al., 2024).
In our proposed method, we leverage the SOTA RL training method, GRPO (Zhihong et al., ), as our basic reinforcement learning method. However, we modify the reward function to make the GRPO have a more stable process and one that is suitable for manipulating the browser.
3. Method
In this section, we present a detailed description of the proposed method. Firstly, we introduce an automatic process for generating verifiable tasks to manipulate web browsers and other computer applications. Subsequently, we elaborate on the sample purification method leveraging LLM-as-judge; this approach ensures the data exhibits a favorable difficulty distribution, which facilitates effective model learning. Following this, we provide a detailed explanation of the proposed reinforcement learning (RL) method. Specifically, we first review the GRPO algorithm, then describe how we adapt it for training our model through modifications to the learning process and reward function. Next, we introduce the Grounding-Based Generalization Enhancement module, which leverages the model’s grounding capability to enhance its planning performance—thereby enabling the model to effectively manipulate target applications.
3.1. Data Engine for Self-evolution
One of the key factors driving model evolution is the ability to scale training data, as this directly enhances the model’s capacity for exploration. However, collecting human-like computer operation data—data that mimics real human interaction with computers—is highly complex, leading to insufficient volumes that hinder effective data scaling. To address this limitation, we propose a novel approach for generating a trajectory dataset of operations on diverse computer software. Specifically, this approach comprises two core components: (1) Generation of closed-loop verifiable tasks; (2) Generation and Evaluation of Trajectory Extraction.
3.1.1. Generation of closed-loop verifiable tasks
To enable the agent to better handle complex multimodal tasks and dynamic environment interactions, we design a closed-loop task data generation pipeline, as illustrated in Figure 2.
@@ 3.1.1. Generation of closed-loop verifiable tasks
Firstly, the proposed pipeline will use a Task Agent to produce many task instructions. The task agent will use some actual tasks of a specific software as the few-shot examples to evoke the agent’s in-context learning of the target sofware. During the generation process, the task agent will periodically check whether the generated tasks can be executed and verified and whether there are some similar tasks have been generated. Meanwhile, the agent will produce some guidelines of how to complete the task. Secondly, the code generation agent will take the generated task instruction and the guideline as inputs to generate a batch of execution and verification program. For execution program, it’s target is to achieve tasks’ goals described in instructions by using the Python. And for the verification program, it judges whether the execution programs’ output matches the tasks’ requirement. In the end, we will automatically run pairs of execution program and verification to select the correct verification program. So in this way, we can get the verifiable tasks.
Figure 2. Illustration of the closed-loop task data generation pipeline. The pipeline consists of two core agents: (1) Task Agent: Generates task instructions (using few-shot examples of real software tasks) and checks executability/duplication; (2) Coding Agent: Takes task instructions and guidelines as input to synthesize Python-based execution programs (for task completion) and verification programs (for validating task success). Only tasks passing automatic execution and verification are retained.
Figure 3. Illustration of the data generation and multi-stage trajectory filtering strategy. This figure illustrates the data generation and multi-stage trajectory filtering strategy. Beginning with task instructions, execution programs, and verification programs, our SEA generates multiple trajectories. These trajectories subsequently go through rule-based selection and filtering by a step filter model. Eventually, the filtered outcomes are utilized to periodically replace and update our SEA for performance refinement.
3.1.2. Generation and Assessment for Trajectory Extraction
@@
To enable the agent to learn the most relevant and effective action sequences, we propose a novel multi-stage trajectory filtering strategy, as illustrated in Figure 3. After getting the task instruction and verification program, we will firstly use our cold-start model, UI-TARS-1.5-7B (Yujia et al., 2025), to perform multiple rounds of inference and obtain diverse task trajectories. And the verification program is then used to select only those trajectories that lead to successful task completion. For each task, we prioritize trajectories with the minimal number of steps based on step-length information. Additionally, to further ensure the quality of each step within the trajectory, we train a step filtering model to filter out redundant or erroneous steps, thereby yielding optimal training trajectories. And the step model will be introduced in the next section. It is notable that during the training process, we will gradually replace the cold-start model with our trained model to generate trajectories for later training stages. In this way, we can make our model evolve through the whole training phase.
3.1.3. Filtering data with step model